# Analyzing LLMs’ Knowledge Boundary Cognition Across Languages Through the Lens of Internal Representations

### v1

@@ Abstract
While understanding the knowledge boundaries of LLMs is crucial to prevent hallucination, research on knowledge boundaries of LLMs has predominantly focused on English. In this work, we present the first study to analyze how LLMs recognize knowledge boundaries across different languages by probing their internal representations when processing known and unknown questions in multiple languages. Our empirical studies reveal three key findings: 1) LLMs’ perceptions of knowledge boundaries are encoded in the middle to middle-upper layers across different languages. 2) Language differences in knowledge boundary perception follow a linear structure, which motivates our proposal of a training-free alignment method that effectively transfers knowledge boundary perception ability across languages, thereby helping reduce hallucination risk in low-resource languages; 3) Fine-tuning on bilingual question pair translation further enhances LLMs’ recognition of knowledge boundaries across languages. Given the absence of standard testbeds for cross-lingual knowledge boundary analysis, we construct a multilingual evaluation suite comprising three representative types of knowledge boundary data. Our code and datasets are publicly available at https://github.com/DAMO-NLP-SG/LLM-Multilingual-Knowledge-Boundaries.
@@
Chenghao Xiao1,2, [^2] Work done during internship at Alibaba DAMO Academy.   Hou Pong Chan1, [^3] Corresponding authors.  Hao Zhang1, [^3] Corresponding authors.   Mahani Aljunied1 Lidong Bing1   Noura Al Moubayed2   Yu Rong1 1 DAMO Academy, Alibaba Group 2 Department of Computer Science, Durham University {xiaochenghao.xch,houpong.chan,hz.hhea2e}@alibaba-inc.com
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has revolutionized their capacity to store and utilize knowledge across languages Dubey et al. (2024); Qwen et al. (2025). Understanding the knowledge boundaries of Large Language Models (LLMs) is critical, as LLMs tend to hallucinate when attempting to answer questions beyond their knowledge. Yet, research on knowledge boundaries of LLMs has predominantly focused on English Azaria and Mitchell (2023); Marks and Tegmark (2024); Li et al. (2024). Misaligned knowledge boundaries between languages can lead to inconsistent and unsafe outputs in cross-lingual applications. Therefore, it is crucial to determine whether the knowledge boundary perceptions of LLMs observed in English can be similarly identified in other languages or transferred across them.
@@
Motivated by the above observations, we further explore whether a specific structure exists in the geometry of multilingual knowledge boundary representations, allowing training-free alignment methods to transfer abilities between languages. Experiments (§5) show that language difference for knowledge boundary is encoded in a linear structure, and training-free alignment methods like as mean-shifting and linear projection can effectively transfer the knowledge boundary perception across languages. Notably, linear projection largely closes the performance gap between in-distribution (ID) and out-of-distribution (OOD) language representations, demonstrating improved cross-lingual transferability. We also find that projecting high-resource language representations onto low-resource subspaces removes extraneous noises encoded by the inherent high dimensionality of high-resource languages. As a result, compared to the low-resource language representation themselves, probes trained on low-resource languages achieve better performance on these projected high-resource language representations than on low-resource language representation themselves, revealing a “weak-to-strong generalization” pattern.
@@
To address the current lack of standard testbeds for evaluating the generalization of knowledge boundary cognition across languages, we constructed three types of multilingual knowledge boundary datasets including: 1) questions with true/false premises; 2) entity-centric answerable/unanswerable questions; and 3) true/false statements. For true/false-premise questions, we first contribute an augmented version of FreshQA Vu et al. (2024) by flipping the premise of each question, and creating its corresponding multilingual version. For entity-centric answerable/unanswerable questions, we create a multilingual QA dataset using an entity-swapping framework, where the content is written in authentic language for each target language, based on a diverse array of real and fictional entities. We also translate the widely-used true/false statement dataset in Azaria and Mitchell (2023) into 8 languages.
@@ 2.1 Knowledge Boundaries of LLMs
While knowledge boundaries are crucial for mitigating hallucinations and unsafe generation, the concept remains underdefined Li et al. (2024). From a generative perspective, Yin et al. (2024) classify LLM’s knowledge into prompt-agnostic knowledge, prompt-sensitive knowledge, and unknown knowledge. Complementarily, representations-based studies focus on whether LLMs’ knowledge boundary perception is reflected in their internal states, through probing LLM’s representations with true/false statements and analyzing their geometric patterns Azaria and Mitchell (2023); Marks and Tegmark (2024); Bürger et al. (2024). However, most work in this line is limited in true/false statements, and no work has systematically investigated multilingual knowledge boundaries, except Bürger et al. (2024) briefly showed that probes trained on English show certain generalization to German statements. In this work, we show that most inter-language generalization gap can be closed with training-free subspace alignment.
@@ 2.2 Multilingualism in LLMs
Multilingual large language models increasingly support diverse languages through larger-scale training Qwen et al. (2025); Dubey et al. (2024) and targeted efforts towards multilinguality BigScience Workshop (2023); Üstün et al. (2024); Zhang et al. (2024c). Studies identify language-specific neurons in multilingual large language models and steer LLMs behaviors by perturbing such neurons Tang et al. (2024); Zhao et al. (2024), revealing sparse language-specific neurons and “anchor language” processing in middle layers Zhao et al. (2024). Mu et al. (2024) link parallel multilingual inputs during inference to inhibited neurons and precise neuron activation. Cross-lingual improvement patterns emerge during fine-tuning: Zhang et al. (2024b) find that question translation fine-tuning boosts multilingual performance, even for languages that are not directly fine-tuned on. In this work, we provide more concrete evidence to this method from a representation perspective.
@@ 3 Datasets
Our FreshQAParallel dataset extends FreshQA Vu et al. (2024) by creating parallel true/false premise questions. Human annotators are asked to manually invert the premise type of each question, e.g., converting “When did Google release ChatGPT?” (false-premise) to “When did OpenAI release ChatGPT?” (true-premise). We highlight the difficulty of converting true-premise question to false-premise due to larger search space, and provide more details of annotation in Appendix F. We further apply GPT-4o to translate the dataset into 7 languages, which are then quality-checked by linguists in the team.
@@
We propose the SeaRefuse Benchmark, which is composed of unanswerable questions about non-existing entities and answerable questions in English, Chinese, Indonesian, Thai, and Vietnamese. We devise an entity-swapping approach and a generation-based approach to construct the questions about fake entities. The entity swapping approach constructs unanswerable questions by swapping the named entities in questions from open-source QA datasets into non-existing entities. The questions form SeaRefuse-H training/test set. Each question in the test set is verified by linguists. The generation-based approach relies on GPT-4o to synthesize unanswerable questions, resulting in the SeaRefuse-G test set. More details are in Appendix G.
@@ 4 Knowledge Boundary Probing
RQ1.1: Do probes transfer cross-lingually, and what’s the pattern across layers, languages and models?
@@
Figure 3 visualizes the transferability matrices of probing model on Qwen2.5-7B’s 3rd and 19th layers, which exhibit the worst and best average transferability across langauges, respectively. Figure 2 reveals a significant gap between in-distribution (blue line) and OOD (orange line) performance, which is most pronounced in bottom layers. In these layers, language models begin incorporating language-specific static embeddings and processing within corresponding language subspaces before converging to a unified knowledge representation space shared across languages in the middle and mid-upper layers.
@@
Above findings hold across model families (Llama, Qwen), sizes (7B to 72B), and variants (base/instruct). We refer to Appendix for full results across models. Representations show subpar cross-lingual transferability in bottom layers. The transferability peaks in middle (Llama) and mid-upper layers (Qwen), then diverges again in final layers. Our results offer more concrete and direct evidence for the 3-phase hypothesis in Zhao et al. (2024), where language models first process input in the source language, then “think” in English or other anchor languages (such as English and Chinese for Qwen), and finally switch back to the source language for generation.
@@
Table 2: Performance comparison between model without and with hints suggesting the question may be false-premise (Baseline vs. FP-Hinted).
@@ 5.2 Subspace Projections
Figure 4: We project the 19th layer of Qwen2.5-7B onto a linear subspace where x-axis encodes languages, y-axis encodes topic-aware true/falsity, and z-axis encodes binary truth/falsity. We visualize the same scatter plot with three different ground-truth label sets annotated respectively.
@@
This difference vector $\Delta\boldsymbol{\mu}$ represents the translation needed to align the mean of the OOD subspace with that of the in-distribution subspace, which is then applied to the OOD test set $\mathbf{X}_{\text{ood}}^{\text{test}}\in\mathbb{R}^{p\times d}$ to obtain the shifted test set $\mathbf{X}_{\text{shifted}}^{\text{test}}$: $\mathbf{X}_{\text{shifted}}^{\text{test}}=\mathbf{X}_{\text{ood}}^{\text{test}}+\Delta\boldsymbol{\mu}.$ Appendix J shows that language means can also be approximated with non-parallel corpus.
@@
Once $\mathbf{W}$ is estimated, it can be used to project the out-of-distribution test set $\mathbf{X}_{\text{ood}}^{\text{test}}\in\mathbb{R}^{p\times d}$ into the in-distribution subspace, resulting in the shifted test set $\mathbf{X}_{\text{shifted}}^{\text{test}}$: $\mathbf{X}_{\text{shifted}}^{\text{test}}=\mathbf{X}_{\text{ood}}^{\text{test}}\mathbf{W}$. For scalability, inter-language projection matrices can be precomputed by sampling $\mathbf{X}{\text{ood}}$ and $\mathbf{X}{\text{in}}$ from larger parallel corpus $\mathbb{C}_{\text{ood}}$ and $\mathbb{C}_{\text{in}}$.
@@
Effective Dimensionality Participation Ratio
Original Projected Original Projected
Khmer 103 97 15.93 18.07
English 116 87 26.26 19.29
@@ 5.3 Findings
Table 3 contrasts the results in English (high-resource) and Khmer (low-resource). English’s higher intrinsic effective dimensionality suggests it encodes more features, potentially language-specific and task-irrelevant details. In contrast, Khmer’s subspace is more compact and task-focused, likely due to limited training data that force LLMs to prioritize generalizable features during pretraining. However, projecting English onto Khmer’s subspace reduces its effective dimensionality ($116\to 87$), stripping away noise or redundant features. English’s higher PR indicates diffuse variance across many directions, whereas Khmer’s lower PR reflects tighter feature alignment with the tasks. English $\to$ Khmer subspace reduces PR ($26.26\to 19.29$), aligning closer to Khmer’s compactness. Khmer → English subspace increases PR ($15.93\to 18.07$), indicating a loss of focus. Thus, low-resource subspaces act as inductive biases for cross-lingual tasks by filtering noise and preserving core semantics, while high-resource subspaces risk carrying superfluous details that degrade transfer.
@@ 6 Training-based Enhancement
We fine-tune Llama-3.1 and Qwen2.5 (from 7B to 72B) via SFT on question translation pairs (e.g., English-Lao), which are constructed using the English subset of SeaRefuse, translated by GPT-4o and verified by linguists. We measure the performance post-SFT using the same protocols in §4 and §5. More details are presented in Appendix E.
@@ Appendix D Pattern 2 Full Visualization
Appendix E Implementation Details of SFT-based Method
We conduct experiments on all model sizes with both Lora and Full-parameter fine-tuning. For Lora, we use a rank 32 and alpha 64. We use a learning rate of $51e-$ 5 for Full-parameter fine-tuning and $41e-$ 4 for Lora, a global batch size of $512$ for both modes through different per device batch size and gradient accumulation steps for models of different sizes. Qwen2.5-72B is trained on 4 nodes of 32 A800 80G GPUs in total with DeepSpeed ZeRO-3 for full-parameter training.
Appendix F Details of Data Construction of FreshQAParallel
Although flipping a false-premise question to true-premise could be more simple by referring to search engines or reasoning based on the groundtruth answers about the original questions in FreshQA, flipping a true-premise question to false-premise is more difficult, given the larger search space (many ways to make it false-premise). We note that, the flipped question must be talking about the same topic with the original question; without getting too creative (e.g., GPT-4o would frequently fail to do this task as it generates too fictional entities such as transforming QS World Ranking to Hogwarts Ranking); and not grounded in the futuristic property relative to model’s training cutoff time (e.g., Who is the US president in 2024 is not deemed a false-premise question to a model trained in 2023, although it’s unanswerable without retrieval-augmented generation; However “Who was the US president in 2035” is false-premise, which assumes a future date is in the past).
@@ Appendix F Details of Data Construction of FreshQAParallel
Example 1.b. “How old is Leonardo DiCaprio’s second kid?” This is a valid false-premise question, because it has assumed a false fact (that the entity has 2 or more kids).
Similarly, Example 2.a. “How many championships has Chris Paul won?” (not a valid false-premise question because it doesn’t assume anything; the answer can be 0.) Example 2.b. “When did Chris Paul win the second championship?” (a valid false-premise question because it has assumed a false fact.)
@@
Appendix G Details of Data Construction of SeaRefuse

### v2

@@ Abstract
While understanding the knowledge boundaries of LLMs is crucial to prevent hallucination, research on the knowledge boundaries of LLMs has predominantly focused on English. In this work, we present the first study to analyze how LLMs recognize knowledge boundaries across different languages by probing their internal representations when processing known and unknown questions in multiple languages. Our empirical studies reveal three key findings: 1) LLMs’ perceptions of knowledge boundaries are encoded in the middle to middle-upper layers across different languages. 2) Language differences in knowledge boundary perception follow a linear structure, which motivates our proposal of a training-free alignment method that effectively transfers knowledge boundary perception ability across languages, thereby helping reduce hallucination risk in low-resource languages; 3) Fine-tuning on bilingual question pair translation further enhances LLMs’ recognition of knowledge boundaries across languages. Given the absence of standard testbeds for cross-lingual knowledge boundary analysis, we construct a multilingual evaluation suite comprising three representative types of knowledge boundary data. Our code and datasets are publicly available at https://github.com/DAMO-NLP-SG/LLM-Multilingual-Knowledge-Boundaries.
@@
Chenghao Xiao1,2, [^2] Work done during internship at Alibaba DAMO Academy.   Hou Pong Chan1, [^3] Corresponding authors.  Hao Zhang1, [^3] Corresponding authors.   Mahani Aljunied1 Lidong Bing1   Noura Al Moubayed2   Yu Rong1,3 1 DAMO Academy, Alibaba Group 2 Department of Computer Science, Durham University 3 Hupan Lab, Hangzhou, China {xiaochenghao.xch,houpong.chan,hz.hhea2e}@alibaba-inc.com
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has revolutionized their capacity to store and utilize knowledge across languages Dubey et al. (2024); Qwen et al. (2025). Understanding the knowledge boundaries of Large Language Models (LLMs) is critical, as LLMs tend to hallucinate when attempting to answer questions beyond their knowledge. Yet, research on knowledge boundaries of LLMs has predominantly focused on English Azaria and Mitchell (2023); Marks and Tegmark (2024); Li et al. (2024); DBLP:conf/emnlp/CheangCW0LS0C23. Misaligned knowledge boundaries between languages can lead to inconsistent and unsafe outputs in cross-lingual applications. Therefore, it is crucial to determine whether the knowledge boundary perceptions of LLMs observed in English can be similarly identified in other languages or transferred across them.
@@
Motivated by the above observations, we further explore whether a specific structure exists in the geometry of multilingual knowledge boundary representations, allowing training-free alignment methods to transfer abilities between languages. Experiments (§5) show that language difference for knowledge boundary is encoded in a linear structure, and training-free alignment methods, such as mean-shifting and linear projection, can effectively transfer the knowledge boundary perception across languages. Notably, linear projection largely closes the performance gap between in-distribution (ID) and out-of-distribution (OOD) language representations, demonstrating improved cross-lingual transferability. We also find that projecting high-resource language representations onto low-resource subspaces removes extraneous noise encoded by the inherent high dimensionality of high-resource languages. As a result, compared to the low-resource language representation themselves, probes trained on low-resource languages achieve better performance on these projected high-resource language representations than on low-resource language representation themselves, revealing a “weak-to-strong generalization” pattern.
@@
To address the current lack of standard testbeds for evaluating the generalization of knowledge boundary cognition across languages, we constructed three types of multilingual knowledge boundary datasets, including: 1) questions with true/false premises; 2) entity-centric answerable/unanswerable questions; and 3) true/false statements. For true/false-premise questions, we first contribute an augmented version of FreshQA Vu et al. (2024) by flipping the premise of each question, and creating its corresponding multilingual version. For entity-centric answerable/unanswerable questions, we create a multilingual QA dataset using an entity-swapping framework, where the content is written in authentic language for each target language, based on a diverse array of real and fictional entities. We also translate the widely used true/false statement dataset in Azaria and Mitchell (2023) into 8 languages.
@@ 2.1 Knowledge Boundaries of LLMs
While knowledge boundaries are crucial for mitigating hallucinations and unsafe generation, the concept remains underdefined Li et al. (2024). From a generative perspective, Yin et al. (2024) classify LLM’s knowledge into prompt-agnostic knowledge, prompt-sensitive knowledge, and unknown knowledge. Complementarily, representations-based studies focus on whether LLMs’ knowledge boundary perception is reflected in their internal states, through probing LLM’s representations with true/false statements and analyzing their geometric patterns Azaria and Mitchell (2023); Marks and Tegmark (2024); Bürger et al. (2024). However, most work in this line is limited in true/false statements, and no work has systematically investigated multilingual knowledge boundaries, except Bürger et al. (2024) briefly showed that probes trained on English show certain generalization to German statements. In this work, we show that most inter-language generalization gaps can be closed with training-free subspace alignment.
@@ 2.2 Multilingualism in LLMs
Multilingual large language models increasingly support diverse languages through larger-scale training Qwen et al. (2025); Dubey et al. (2024) and targeted efforts towards multilinguality BigScience Workshop (2023); Üstün et al. (2024); Zhang et al. (2024c). Studies identify language-specific neurons in multilingual large language models and steer LLMs’ behaviors by perturbing such neurons Tang et al. (2024); Zhao et al. (2024), revealing sparse language-specific neurons and “anchor language” processing in middle layers Zhao et al. (2024). Mu et al. (2024) link parallel multilingual inputs during inference to inhibit neurons and precise neuron activation. Cross-lingual improvement patterns emerge during fine-tuning: Zhang et al. (2024b) find that question translation fine-tuning boosts multilingual performance, even for languages that are not directly fine-tuned on. In this work, we provide more concrete evidence to this method from a representation perspective.
@@ 3 Datasets
Our FreshQAParallel dataset extends FreshQA Vu et al. (2024) by creating parallel true/false premise questions. Human annotators are asked to manually invert the premise type of each question, e.g., converting “When did Google release ChatGPT?” (false-premise) to “When did OpenAI release ChatGPT?” (true-premise). We highlight the difficulty of converting true-premise questions to false-premise due to larger search space, and provide more details of the annotation in Appendix E. We further apply GPT-4o to translate the dataset into 7 languages, which are then quality-checked by linguists in the team.
@@
We propose the SeaRefuse Benchmark, which is composed of unanswerable questions about non-existing entities and answerable questions in English, Chinese, Indonesian, Thai, and Vietnamese. We devise an entity-swapping approach and a generation-based approach to construct questions about fake entities. The entity swapping approach constructs unanswerable questions by swapping the named entities in questions from open-source QA datasets into non-existent entities. The questions form SeaRefuse-H training/test set. Each question in the test set is verified by linguists. The generation-based approach relies on GPT-4o to synthesize unanswerable questions, resulting in the SeaRefuse-G test set. More details are in Appendix F.
@@ 4 Knowledge Boundary Probing
RQ1.1: Do probes transfer cross-lingually, and what’s the pattern across layers, languages, and models?
@@
Figure 3 visualizes the transferability matrices of probing model on Qwen2.5-7B’s 3rd and 19th layers, which exhibit the worst and best average transferability across languages, respectively. Figure 2 reveals a significant gap between in-distribution (blue line) and OOD (orange line) performance, which is most pronounced in the bottom layers. In these layers, language models begin incorporating language-specific static embeddings and processing within corresponding language subspaces before converging to a unified knowledge representation space shared across languages in the middle and mid-upper layers.
@@
The above findings hold across model families (Llama, Qwen), sizes (7B to 72B), and variants (base/instruct). We refer to the Appendix for full results across models. Representations show subpar cross-lingual transferability in the bottom layers. The transferability peaks in middle (Llama) and mid-upper layers (Qwen), then diverges again in final layers.
Our results offer more concrete and direct evidence for the 3-phase hypothesis in Zhao et al. (2024), where language models first process input in the source language, then “think” in English or other anchor languages (such as English and Chinese for Qwen), and finally switch back to the source language for generation.
@@
Table 2: Performance comparison between models without and with hints suggesting the question may be false-premise (Baseline vs. FP-Hinted).
@@ 5.2 Subspace Projections
Figure 4: We project the 19th layer of Qwen2.5-7B onto a linear subspace where the x-axis encodes languages, the y-axis encodes topic-aware true/falsity, and the z-axis encodes binary truth/falsity. We visualize the same scatter plot with three different ground-truth label sets annotated respectively.
@@
This difference vector $\Delta\boldsymbol{\mu}$ represents the translation needed to align the mean of the OOD subspace with that of the in-distribution subspace, which is then applied to the OOD test set $\mathbf{X}_{\text{ood}}^{\text{test}}\in\mathbb{R}^{p\times d}$ to obtain the shifted test set $\mathbf{X}_{\text{shifted}}^{\text{test}}$: $\mathbf{X}_{\text{shifted}}^{\text{test}}=\mathbf{X}_{\text{ood}}^{\text{test}}+\Delta\boldsymbol{\mu}.$ Appendix I shows that language means can also be approximated with non-parallel corpus.
@@
Once $\mathbf{W}$ is estimated, it can be used to project the out-of-distribution test set $\mathbf{X}_{\text{ood}}^{\text{test}}\in\mathbb{R}^{p\times d}$ into the in-distribution subspace, resulting in the shifted test set $\mathbf{X}_{\text{shifted}}^{\text{test}}$: $\mathbf{X}_{\text{shifted}}^{\text{test}}=\mathbf{X}_{\text{ood}}^{\text{test}}\mathbf{W}$. For scalability, inter-language projection matrices can be precomputed by sampling $\mathbf{X}{\text{ood}}$ and $\mathbf{X}{\text{in}}$ from larger parallel corpora $\mathbb{C}_{\text{ood}}$ and $\mathbb{C}_{\text{in}}$.
@@ 5.3 Findings
Effective Dimensionality Participation Ratio
Original Projected Original Projected
Khmer 103 97 15.93 18.07
English 116 87 26.26 19.29
Table 3 compares the results in English (high-resource) and Khmer (low-resource). English’s higher intrinsic effective dimensionality suggests it encodes more features, potentially language-specific and task-irrelevant details. In contrast, Khmer’s subspace is more compact and task-focused, likely due to limited training data that forces LLMs to prioritize generalizable features during pretraining. However, projecting English onto Khmer’s subspace reduces its effective dimensionality ($116\to 87$), stripping away noise or redundant features. English’s higher PR indicates diffuse variance across many directions, whereas Khmer’s lower PR reflects tighter feature alignment with the tasks. English $\to$ Khmer subspace reduces PR ($26.26\to 19.29$), aligning closer to Khmer’s compactness. Khmer → English subspace increases PR ($15.93\to 18.07$), indicating a loss of focus. Thus, low-resource subspaces act as inductive biases for cross-lingual tasks by filtering noise and preserving core semantics, while high-resource subspaces risk carrying superfluous details that degrade transfer.
@@ 6 Training-based Enhancement
We fine-tune Llama-3.1 and Qwen2.5 (from 7B to 72B) via SFT on question translation pairs (e.g., English-Lao), which are constructed using the English subset of SeaRefuse, translated by GPT-4o and verified by linguists. We measure the performance post-SFT using the same protocols in §4 and §5.
We conduct experiments on all model sizes with both Full-parameter fine-tuning and LoRA. For LoRA, we use a rank 32 and alpha 64. We use a learning rate of $51e-$ 5 for full fine-tuning and $41e-$ 4 for LoRA. A global batch size of $512$ is maintained for all settings through different gradient accumulation setups. Qwen2.5-72B is trained on 4 nodes of 32 A800 80G GPUs in total with DeepSpeed ZeRO-3 for full-parameter training.
@@ Appendix D Pattern 2 Full Visualization
Appendix E Details of Data Construction of FreshQAParallel
Although flipping a false-premise question to true-premise could be simpler by referring to search engines or reasoning based on the groundtruth answers about the original questions in FreshQA, flipping a true-premise question to false-premise is more difficult, given the larger search space (many ways to make it false-premise). We note that, the flipped question must be talking about the same topic with the original question; without getting too creative (e.g., GPT-4o would frequently fail to do this task as it generates too fictional entities such as transforming QS World Ranking to Hogwarts Ranking); and not grounded in the futuristic property relative to model’s training cutoff time (e.g., Who is the US president in 2024 is not deemed a false-premise question to a model trained in 2023, although it’s unanswerable without retrieval-augmented generation; However “Who was the US president in 2035” is false-premise, which assumes a future date is in the past).
@@ Appendix E Details of Data Construction of FreshQAParallel
Example 1.b. “How old is Leonardo DiCaprio’s second kid?” This is a valid false-premise question because it has assumed a false fact (that the entity has 2 or more kids).
Similarly, Example 2.a. “How many championships has Chris Paul won?” (not a valid false-premise question because it doesn’t assume anything; the answer can be 0). Example 2.b. “When did Chris Paul win the second championship?” (a valid false-premise question because it has assumed a false fact.)
@@
Appendix F Details of Data Construction of SeaRefuse